Margaret Moore is saving to buy a house in five years. She plans to put 20 percent down at that time, and she believes that she will need $32,000 for the down payment. If Margaret can invest in a fund that pays 6.20 percent annual interest, compounded quarterly, how much will she have to invest today to have enough money for the down payment? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest penny.)
Answer #1

Amount required after 5 years = $32,000

Number of quarters = 5*4 = 20

Quarterly Rate = 6.20%/4 = 1.55%

Amount required to be invested today is equal to present value of future amount

= 32,000*PVF(1.55%, 20 periods)

= 32,000*0.735193

= $23,526.18